# Localizing Before Answering: A Hallucination Evaluation Benchmark for Grounded Medical Multimodal LLMs

### v3

@@ 4.1 Grounded Text Generation
We use the projected hidden state $h_{\text{seg}}$ from the <SEG> token as the query to a the segmentation model. This can be formulated as $\hat{M}=Seg(x_{\text{img}},h_{\text{seg}})$. Our LLM is trained via the text objectives. Similar to other LLMs Touvron et al. (2023); OpenAI et al. (2023), we formulate the text generation loss as the cross entropy loss between the generated text and the ground truth
@@
Finally, the overall objective can be seen as the weighted sum of the two losses
@@ 4.2 Attention Highlighter via Self-Prompting
For attention reweighting, patches of interest $x_{hl}$ are added extra weight before feeding through the softmax layer so that those regions get more focus on during answer generation. Denote $h_{j,i}$ the original value of the attention logits of patch j to patch $i$, and $\beta$ the hyperparameter of added weight, then the new attention weight after reweighting $\hat{h}_{j,i}$ becomes
@@
Here, $m_{i}=1$, indicating token $i$ is highlighted, and $m_{i}=0$ otherwise. We apply this function to the visual backbone of M-LLM where the patches of interests are interpolated from the mask. Applying the reweighted attention $\tilde{A}$, we obtain the highlighted image tokens $\tilde{x}_{\text{img}}$:
@@
To further alleviate shortcut learning, we perform contrastive decoding Zhang et al. (2024b), which contrasts the model decisions after and before highlighting regions of interests. Particularly, we compute the difference between the token probability after highlighting $p_{hl}$ and the token probability before highlighting $p_{bh}$ to obtain the decoded probability $p$:
@@
where $\alpha$ is a hyperparameter controlling the degree of grounding. Applying contrastive decoding minimizes the effect of other tokens to the model output, thus preventing hallucination.
Discussion. Although Prompt Highlighter Zhang et al. (2024b) also emphasizes the “highlighting” of regions of interest to reduce hallucination, ours differ in two folds. First, Prompt Highlighter requires human prompting for attention highlighting, which is expensive in medical image analysis. In contrast, ours automatically localizes regions of interests, thus obviating the need for medical professionals to manually localize regions of interest. Second, Prompt Highlighter only enhances attention on the language decoder, our methodology reweighs attention scores on the visual backbone. The deep intervention on visual backbone is more effective in mitigating the visual shortcut learning.
@@ 5 Experimental Results
Module Backbone LobA LobA
w/o self-prompt (Ours)
@@
Benchmarks. We conduct a thorough and systemic evaluation of the most popular Medical Large Multimodal Models on our large-scale HEAL-MedVQA benchmark. Specifically, we utilize 7 state-of-the-art multi-modal LLMs for our evaluation protocol, 3 of which are proprietary models: GPT-4o, GPT-4 Vision OpenAI et al. (2023) and Gemini 1.5 Team et al. (2023) and 4 are open source M-LLMs of the medical domain: CheXagent Chen et al. (2024), LLaVA-Med Li et al. (2024), BioMedGPT Zhang et al. (2023a) and Med-Flamingo Alayrac et al. (2022). We also assess the capabilities of state-of-the-art adversarial VQA methods, including VCD Leng et al. (2024), Prompt Highlighter Zhang et al. (2024b), and CRG Wan et al. (2025).
@@
Evaluation Metrics. To assess the accuracy of the multi-modal LLMs, we utilize a powerful LLM, LLaMa 3.1-8B specifically, capable of extracting information from the M-LLM’s output, and categorizing them into a set of disease labels. For example, for an open-ended question, if the model returns the answer “The heart suffers from pneumonia, pulmonary fibrosis and nodule/mass” then the extracted labels by the LLM will be pneumonia, pulmonary fibrosis, nodule/mass. We report the multi-label precision, recall, and micro F1 score as the main accuracy metrics. For hallucination sensitivity analysis, we adopt our proposed evaluation metrics, Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT) scores, as discussed in Sec. 3.1.
@@
Experimental Results and Discussion. VQA Accuracy. Table LABEL:tab:performance_comparison reports the quantitative accuracy of model’s answers on MIMIC and VinDr datasets. Proprietary models generally achieve sub-optimal accuracy, with the accuracy being less than 50% for most models on binary Yes/No questions and mostly less than 10% for open-ended questions. This shows the challenging nature of our HEAL-MedVQA benchmark, requiring the model to be robust against visual and textual shortcut learning. Our proposed framework consistently outperforms recent advances in M-LLMs and adversarial VQA techniques even after fine-tuning them. Notably, the proposed LoBA outperforms the state-of-the-art baseline Prompt Highlighter Zhang et al. (2024b) by 3.19% and 3.63% on Open-ended questions for VinDr and MIMIC-CXR, respectively. This shows the effectiveness of grounding on visual evidence before answering as in our proposal.
@@
Textual Perturbation Test. Table 2 reports the sensitivity of each model to textual language bias, measuring the percentage of changed True Positive answers when either switching an anatomy or a disease. The higher the percentage is, the more robust a model is from textual shortcut bias. Our proposed framework achieves the highest sensitivity measure overall, with the ratio of changed answers drastically increasing by up to 3.89% compared to CRG Wan et al. (2025). Enhancing attention to the queried regions, as in our proposal, helps alleviate the textual and visual shortcut learning. Lastly, fine-tuned open-source models are generally more robust than closed-source models like GPT-4-O, showing that fine-tuning LLMs on our adversarial data remedies their hallucination problems.
@@
Visual Perturbation Test. Table 3 reports the visual sensitive score, with the score defined as the number of changed answers when the original image is overlaid with a healthy localized anatomy. Most multi-modal LLMs in our benchmark shift their answer accordingly around 40-70% of the time in our experiments, indicating the big influence of the area of interest on a model’s output. Our method is the most sensitive to the change in the localized area, increasing the number of visually dependent answers by 1.94% compared to Prompt Highlighter. Zhang et al. (2024b)
@@
Ablation studies. Table 4 reports the effects of each component in our LobA framework: the LLaVA Med backbone, Grounded Text Generation module, and the full framework with self-prompting. Applying Grounded Text Generation improves the model’s performance in some categories, showing the effectiveness of localization-aware training.With the self-prompting framework, the accuracy and textual/visual robustness improve greatly, most evidently increasing 4.45% in TPT score, and 2.73% on MIMIC-CXR dataset Yes/No F1 Score.
@@
Qualitative analysis. Fig 6 showcases our LobA and some other LMM’s response on a few qualitative examples, as well as the visual attention map on the input image. We can see that LobA manages to query the attention weights much more intuitively compared to LLaVA-Med or LISA While LLaVA-Med and LISA’s attention are scattered into many irrelevant areas, in the first example LobA’s main attention weights are allocated around the right lower lung area, and the second example is near the left upper lung. Thanks to the grounding generation module, followed by LobA’s self prompting module to automatically re-weight the attention to the area of interest, the model is able to have better visual reasoning capabilities, leading to more accurate answers compared with other methods.
@@ 6 Conclusion
This paper introduces Heal-MedVQA a new large scale Medical Visual Question Answering Benchmark with over 67,000 question-answer pairs, which queries diseases presence at local anatomies, evaluating LMM’s capabilities to localize at grounded visual evidence when answering. In addition, we present the Localize-before-Answer (LobA) framework which trains the LMMs to segment the region of interest and re-adjust their attention for more emphasis on the segmented pathological areas, leading to more reliable answers. Our experimental results showed our framework LobA outperforming state-of-the-art medical LMMs on Heal-MedVQA, proving its robustness and localization capabilities.

### v4

@@ 4.1 Grounded Text Generation
We use the projected hidden state $h_{\text{seg}}$ from the <SEG> token as the query to the segmentation model. This can be formulated as $\hat{M}=Seg(x_{\text{img}},h_{\text{seg}})$. Our LLM is trained via the text objectives. Similar to other LLMs Touvron et al. (2023); OpenAI et al. (2023), we formulate the text generation loss as the cross-entropy loss between the generated text and the ground truth:
@@
Finally, the overall objective can be seen as the weighted sum of the two losses:
@@ 4.2 Attention Highlighter via Self-Prompting
For attention reweighting, patches of interest $x_{hl}$ are added extra weight before feeding through the softmax layer so that those regions get more focus during answer generation. Denote $h_{j,i}$ the original value of the attention logits of patch j to patch $i$, and $\beta$ the hyperparameter of added weight, then the new attention weight after reweighting $\hat{h}_{j,i}$ becomes
@@
Here, $m_{i}=1$, indicating token $i$ is highlighted, and $m_{i}=0$ otherwise. We apply this function to the visual backbone of M-LLM where the patches of interest are interpolated from the mask. Applying the reweighted attention $\tilde{A}$, we obtain the highlighted image tokens $\tilde{x}_{\text{img}}$:
@@
To further alleviate shortcut learning, we perform contrastive decoding Zhang et al. (2024b), which contrasts the model decisions after and before highlighting regions of interest. Particularly, we compute the difference between the token probability after highlighting $p_{hl}$ and probability before highlighting $p_{bh}$ to obtain the decoded probability $p$:
@@
where $\alpha$ is a hyperparameter controlling the degree of grounding. Applying contrastive decoding minimizes the effect of other tokens on the model output, preventing hallucination.
Discussion. Although Prompt Highlighter Zhang et al. (2024b) also emphasizes the “highlighting” of regions of interest to reduce hallucination, ours differs in two ways. First, Prompt Highlighter requires human prompting for attention highlighting, which is expensive in medical image analysis. In contrast, ours automatically localizes regions of interest, thus obviating the need for medical professionals to manually localize regions of interest. Second, while Prompt Highlighter only enhances attention on the language decoder, our methodology reweighs attention scores on the visual backbone. The deep intervention on the visual backbone is more effective in mitigating the visual shortcut learning.
@@ 5 Experimental Results
Module Backbone LobA w/o LobA
self-prompt (Ours)
@@
Benchmarks. We conduct a thorough and systematic evaluation of the most popular Medical Large Multimodal Models on our large-scale HEAL-MedVQA benchmark. Specifically, we utilize 7 state-of-the-art multi-modal LLMs for our evaluation protocol, 3 of which are proprietary models: GPT-4o, GPT-4 Vision OpenAI et al. (2023) and Gemini 1.5 Team et al. (2023) and 4 are open source M-LLMs of the medical domain: CheXagent Chen et al. (2024), LLaVA-Med Li et al. (2024), BioMedGPT Zhang et al. (2023a) and Med-Flamingo Alayrac et al. (2022). We also assess the capabilities of state-of-the-art adversarial VQA methods, including VCD Leng et al. (2024), Prompt Highlighter Zhang et al. (2024b), and CRG Wan et al. (2025).
@@
Evaluation Metrics. To assess the accuracy of the multi-modal LLMs, we utilize LLaMa 3.1-8B to extract information from the M-LLM’s output, and categorize them into a set of disease labels. For example, for an open-ended question, if the model returns the answer “The heart suffers from pneumonia, pulmonary fibrosis and nodule/mass” then the extracted labels by the LLM will be pneumonia, pulmonary fibrosis, nodule/mass. We report the multi-label precision, recall, and micro F1 score as the main accuracy metrics. For hallucination sensitivity analysis, we adopt our proposed evaluation metrics, Textual Perturbation Test (TPT), and Visual Perturbation Test (VPT) scores, as discussed in Sec. 3.1.
@@
Experimental Results and Discussion. VQA Accuracy. As shown in Table LABEL:tab:performance_comparison, proprietary models achieve sub-optimal accuracy. Most models have accuracy less than 50% on binary Yes/No questions and less than 10% for open-ended questions. This shows the challenging nature of our HEAL-MedVQA benchmark, requiring the model to be robust against visual and textual shortcut learning. Our proposed framework consistently outperforms recent advances in M-LLMs and adversarial VQA techniques even after fine-tuning them. Notably, the proposed LoBA outperforms the state-of-the-art Prompt Highlighter Zhang et al. (2024b) by 3.19% and 3.63% on Open-ended questions for VinDr and MIMIC-CXR, respectively. Injecting grounding prior, as in our proposal, boosts the VQA accuracy.
@@
Textual Perturbation Test. Table 2 reports the sensitivity of each model to textual language bias. The higher the percentage is, the more robust a model is against textual shortcut bias. Our proposed framework achieves the highest sensitivity measure overall, with the ratio of changed answers drastically increasing by up to 3.89% compared to CRG Wan et al. (2025). Enforcing visual reasoning via highlighted attention to the queried regions, as in our proposal, alleviates the shortcut learning. Lastly, fine-tuned open-source models are generally more robust than closed-source models like GPT-4-O, showing that fine-tuning LLMs on our adversarial data remedies hallucination.
@@
Visual Perturbation Test. Table 3 reports the visual sensitivity score, with the score defined as the number of changed answers when the original image is overlaid with a healthy localized anatomy. Most multi-modal LLMs in our benchmark shift their answer around 40-70% of the time. Our method is the most sensitive to the change in the localized area, increasing the number of visually dependent answers by 1.94% compared to Prompt Highlighter Zhang et al. (2024b).
@@
Ablation studies. Table 4 reports the effects of each component in our framework: the LLaVA-Med backbone, Grounded Text Generation (GTG) module, and the full framework with self-prompting. Applying GTG improves the model’s performance in some categories, showing the effectiveness of localization-aware training. With the self-prompting framework, the accuracy and textual/visual robustness improve greatly, most evidently increasing 4.45% in TPT score, and 2.73% on MIMIC-CXR dataset Yes/No F1 Score.
@@
Qualitative analysis. Fig. 6 showcases our LobA and some other LMM’s response on some qualitative examples, as well as the visual attention maps on the input images. We can see that LobA manages to query the attention weights much more intuitively compared to LLaVA-Med or LISA. While others’ attentions are scattered into many irrelevant areas, in the first example LobA’s main attention weights are allocated around the right lower lung area, and the second example is near the left upper lung. Thanks to the grounding generation module, followed by LobA’s self-prompting module to automatically re-weight the attention to the area of interest, the model is able to have better visual reasoning capabilities, leading to more accurate answers compared with other methods.
@@ 6 Conclusion
This paper introduces Heal-MedVQA, a new large-scale Medical Visual Question Answering Benchmark with over 67,000 question-answer pairs, which queries diseases at local anatomies, evaluates LMM’s capabilities to localize at grounded visual evidence when answering. In addition, we present the Localize-before-Answer (LobA) framework, which trains the LMMs to segment the region of interest and re-adjust their attention for more emphasis on the segmented pathological areas, leading to more reliable answers. Our experimental results showed that our framework LobA outperformed state-of-the-art medical LMMs on Heal-MedVQA, proving its robustness and localization capabilities.
Acknowledgement
We would like to thank Jiangyu Zhou for assisting with the annotation of chest X-ray images for this project.
Appendix A Computational Efficiency
We present the inference time (seconds) and GPU consumption (MB) of LoBA on a system with a single A100 80GB GPU for 100 samples in comparison with some other medical LMMs in Tab. 5. Our LobA yields computational efficiency comparable to that of existing methods, while achieving superior localization capabilities.
CheXagent Med-Flamingo LLaVA-Med LobA (Ours)
Inf. Time 113.4 126.5 87.0 101.2
GPU Mem 23406 26453 17729 25254
Table 5: Comparison of inference time and GPU consumption
Appendix B Benchmarking Other General LLMs
We report the F1 score of additional proprietary and open-source LMMs (fine-tuned) on the VinDr subset in Tab. 6
Model Gemini-1.5 Claude 3.7 Qwen 2.5 VL InternVL LobA
Pro Sonnet 7B-Instruct 2.5 8B (Ours)
Yes/No 0.583 0.610 0.660 0.654 0.728
Open 0.078 0.062 0.413 0.378 0.542
Table 6: Performance on VinDr for general LMMs
Appendix C Ablation study
We conducted two ablation studies, examining the effect of added attention weights $\beta$ (Tab. 7) and the effect of segmentation quality on the model’s results (Fig 7).
In summary, low weights cause the segmented area to have minimal influence, harming results, while excessive weight (e.g., 4.0) shifts focus away from the question, also degrading performance. For the second ablation study, the quality of the segmentation indeed correlates directly with LoBA’s performance. Overall, our model outperforms LISA baselines when the segmentation Dice score $>0.25$, which accounts for more than $10\%$ of the test set. This shows that our design is already beneficial even when the segmentation module can localize the rough regions to attend to. For samples with Dice $<0.25$, which only account for 9.47% of data, our LoBA yields a slight performance decrease.
$\beta$ Yes/No F1 Open F1
0.5 0.694 0.476
1 0.698 0.510
2 0.720 0.545
3 0.728 0.542
4 0.712 0.501
Table 7: Effect of $\beta$ on performance on VinDr
Figure 7: Yes/No F1 w.r.t segmentation qualities. Data portion (%) on test set is also shown.
$\alpha$ F1 Score
0.0 0.504
0.1 0.509
0.3 0.507
0.5 0.513
1.0 0.517
1.5 0.519
Table 8: F1 Scores for different $\alpha$ values on MIMIC
Appendix D LobA Prompt Template
Figure 8: LobA Instruction Template
Figure 9: GPT and Gemini Models Instruction Template for Yes/No questions
Figure 10: Qualitative Examples from our dataset
Figure 11: GPT and Gemini Models Instruction Template for Open-ended questions
We propose a Grounded Text Generation training paradigm to enable LMM to localize the area of interest with the segmentation token “[SEG]“, before giving its answer for the model to be aware of the localization information.
As such, our model is fine-tuned to follow a strict template as “The location of {relevant areas} is at <SEG>. The answer is …”. To make sure the model’s output comply with this standard, we designed a few-shot instruction template for LobA so that the model can follow the desired template more accurately, displayed in Fig 8
Appendix E Proprietary Models Template
For GPT-4-O and GPT-4-Vision and Gemini-1.5-Flash-8B models, we modify the template of these proprietary models for yes/no questions and open-ended questions separately to allow them to follow our instructions correctly. These 2 templates are displayed by Fig 11 and 9
Appendix F Evaluation Template
Figure 12: Yes/No evaluation prompt template
Figure 13: Open Ended evaluation prompt template
We utilize LLaMA-3 as the LLM backbone to extract the labels from the model’s natural language answers with the following prompt templates, Fig 12 and Fig 13
Appendix G Question example and qualitative results
We present some qualitative examples, as well as some model’s response and their visual attention map in Fig 10.